A compound containing only nitrogen and oxygen is


63.64% N by mass the empirical formula of the


compound is

1
Expert's answer
2022-04-25T14:17:04-0400

Solution:

w(N) = 63.64% or 0.6364

w(O) = 100% − w(N) = 100% − 63.64% = 36.36% or 0.3636


Calculate the mass of each element in 100 g of the compound:

Mass of N = w(N) × Mass of samole = 0.6364 × 100 g = 63.64 g N

Mass of O = w(O) × Mass of sample = 0.3636 × 100 g = 36.36 g O


Calculate the moles of each element in the compound:

Molar mass of nitrogen (N) is 14.0067 g/mol

Molar mass of oxygen (O) is 15.999 g/mol

Therefore

Moles of N = (63.64 g N) × (1 mol N / 14.0067 g N) = 4.5435 mol N

Moles of O = (36.36 g O) × (1 mol O / 15.999 g O) = 2.2726 mol O


Calculate the mole ratio between the elements. Divide the moles of each element by the least number of moles.

N: 4.5435 / 2.2726 = 1.999

O: 2.2726 / 2.2726 = 1.000

Thus, the empirical formula of the compound is N2O


Answer: The empirical formula of the compound is N2O